# BG-GAN: Generative AI Enable Representing Brain Structure-Function Connections for Alzheimer’s Disease

### v3

@@ Abstract
The relationship between brain structure and function is critical for revealing the pathogenesis of brain disease, including Alzheimer’s disease (AD). However, it is a great challenge to map brain structure-function connections due to various reasons. In this work, a bidirectional graph generative adversarial networks (BG-GAN) is proposed to represent brain structure-function connections. Specifically, by designing a module incorporating inner graph convolution network (InnerGCN), the generators of BG-GAN can employ features of direct and indirect brain regions to learn the mapping function between structural domain and functional domain. Besides, a new module named Balancer is designed to counterpoise the optimization between generators and discriminators. By introducing the Balancer into BG-GAN, both the structural generator and functional generator can not only alleviate the issue of mode collapse but also learn complementarity of structural and functional features. Experimental results using ADNI datasets show that the both the generated structure connections and generated function connections can improve the identification accuracy of AD. More importantly, based the proposed model, it is found that the relationship between brain structure and function is not a complete one-to-one correspondence. Brain structure is the basis of brain function. The strong structural connections are almost accompanied by strong functional connections.
@@ I Introduction
Alzheimer’s Disease has placed a huge burden on patients’ families and society as a whole. Symptoms of AD mainly include memory impairment, aphasia, apraxia, cognitive impairment, impairment of visuospatial skills, executive dysfunction, and personality and behavior changes. Although most people experience some degree of memory loss in old age, memory loss is much more severe in people with Alzheimer’s disease than in normal individuals.The term Mild Cognitive Impairment (MCI) was created to differentiate them. Some MCI patients can remain stable for years, while some subjects with AD-like brains change inevitably progress to AD[1]. The prevalence of AD is increasing rapidly due to the fact that the proportion of the population aged 65 and over is growing faster than any other age group worldwide in recent years. Current drugs can relieve symptoms but do not halt disease progression. Learning the relationship between brain structure and function can help doctors diagnose between AD patients and normal people. Moreover, it can also help researchers to figure out the cause of the worsening condition[2, 3, 4].
Recent advancements in artificial intelligence[5, 6, 7, 8] have significantly enhanced the capabilities of neuroimaging analysis, making it more reliable and sensitive than traditional cognitive assessments in detecting AD[9, 10]. Images of different modalities, such as structural Magnetic Resonance Imaging (sMRI), functional Magnetic Resonance Imaging (fMRI), diffusion tensor imaging (DTI) and so on, are well suited to assist physicians in diagnosing patients with the disease [11, 12, 13, 14]. Moreover, with the advantage of high temporal and spatial resolution, each of these tools can identify differences between AD patients and normal individuals well. In the realm of AI-driven diagnostic approaches[15, 16, 17, 18, 19], individual modality is used to study diagnosis between AD and MCI. Some used sMRI brain neuroimages to classify AD and NC [20, 21, 22, 23, 24, 25]. Some used fMRI data [26] and some used DTI data [27]. Due to the limitations of computing power and algorithm at that time[28], they couldn’t utilize the complementary information between multiple modalities’ data[29, 30]. The complex relationships between brain structure and function could not be discovered clearly. The results of diagnosis are not so good as the current methods using multiple modalities images. Now, with the proliferation of artificial intelligence technologies, it is possible to integrate data from multiple imaging modalities.
@@ II Related Works
Zhang et al. [33] recovered the structural connectivity via Multi-GCN from functional connectivity. The recovered result that functional information can predict structural information, revealed that functional information contains all the structural information. Wang et al. [34] proposed an algorithm called CDA and the empirical structural connections (SC) can be reliably predicted based on the direct anatomical relationships, indirect pathways, and module topology interact with one another forming temporally dependent functional connections (FC). Based on the fact that related brain regions with strong functional connectivity can exist without structural connectivity, but the strong functional connectivity can be observed in the brain regions with the strong structural connectivity, there is a many-to-one functional-structural mode.
@@
However, not all researchers agree with the above views. Abdelnour et al [35] tried to use mathematical methods to derive a relationship between SC from DTI and FC from fMRI and found that using the approach via (structural) Laplacian spectral, FC and SC shared eigenvectors and their eigenvalues are exponentially related. Honey et al [36] believed that although resting state functional connectivity is variable, its strength, persistence and spatial statistics are constrained by the large-scale anatomical structure. Messé et al [37] explored the spatial consistency of the relationship between the structure and function of the human brain. The result showed that brain regions can affect the relationship between structural and functional connections. Different division methods will affect the relationship between structure and function. Simon et al [38] discussed the influence of different brain regions on structure and function and proved its effectiveness. Koch et al. [39] examined the correlation of spontaneous fluctuations of brain voxel BOLD signals in the cortical gyrus of the same hemisphere and found that there is a positive correlation between function and structure. Annen et al. [40] explored the relationship of function fluorodeoxyglucose FDG-PET metabolism and structure MRI-diffusion-weighted images (DWI). The fact that the patients’ local metabolism and white matter integrity were significantly reduced, showed that a stronger relationship of function-structure exists in most areas of Default mode networ. Straathof et al. [41] combined multiple articles and draw a conclusion: the structural and functional connection strength of the mammalian brain is positively correlated with the connection strength of the diffuse structure on the macroscopic scale, and positively correlated with the connection strength of the neuron trace structure on the mesoscale.
@@
Based on above analysis, the following simple assumption can be summarized: there is indeed a relationship between structure and function. However, few people study the relationship between structure and function by employing complementary features between brain structure and function.
@@ IV-B BGGAN for mapping domain
Generative adversarial networks [42, 43, 44, 45, 46] commonly consist of two modules: the module generator and the module discriminator. The generator tries to map latent features $z\sim p(z)$ from the noise or from the prior distribution to the distribution of the real domain. The discriminator aims to distinguish between the real data and the fake data. They are trained by the following min-max optimization Equation 1:
@@
In addition, because of the instability of the traditional GAN framework and the unsmoothness of training process, the results of generators are very poor. The generated data are quite different with the source data. In view of above problems, the Balancer modules are add between the source data and the real input of discriminators, which can be seen in section IV-D in details. The loss function of the model can not only calculate the differences in brain region level, but also take into account the overall differences. The details of Balancer modules can be seen in section IV-E.
@@ IV-C InnerGCN for extracting graphic features
Next, we will introduce how to get $U$ in next steps: 1) get the Laplace tensor $L$, where $L(i)=I-D(i)^{-1/2}A(i)D(i)^{-1/2}$; 2) get the $L_{dft}$ from the $L$ in the Fourier form; 3) get the singular values matrix of each dimension of the tensor $[u(i),s(i)]=EVD(L_{dft}(i))$; 4) map the data in the Fourier data back to the original space $U=IDFT(u),\Lambda=IDFT(s)$. $U^{H}$ is the Hermitian matrix of $U$. $U\diamond x$ is the $U$ of Fourier form, and same as convolution filter $g$.
The third-order tensor can be considered as matrix with each element to be a vector. Same with matrix multiplication based on vector multiplication, matrix-matrix interaction is realized through circular convolution based on matrix multiplication. We defined $A,B$ as $A\in\mathbb{R}^{n_{1}\times n_{2}\times 3}$ and $B\in\mathbb{R}^{n_{2}\times n_{3}\times 3}$ and fold1(…), fold2(…), unfold(…) in Eq (6-9):
@@ IV-D Balancer for smoothing training process of BGGAN
A neural network is utilized as Balancer, which is comprised of several convolution networks, a step connection combined the high order information and the low order information from the Balancer. Its loss function contains two parts: an adversarial loss $\mathcal{L}_{B}^{adv}$ and a reconstruction loss $\mathcal{L}_{B}^{recon}$. The function of the adversarial loss is to try fool the discriminator and slow down the discriminators’ learning process. The reconstruction loss is to limit the output similar to real sample and avoid the discriminator learn incorrect distribution too much. The Balancer is trained as following loss:
@@ V Experiments
V-A Datasets
In this subsection, a series of experiments are conducted based on the Alzheimer’s Disease Neuroimaging Initiative (ADNI) datasets[53]. In our experiments, we used 80% subjects as training dataset, used 20% subjects as testing dataset. All experimental results are based on the testing dataset. The details of the datasets used in this paper are shown in Tab.I.
@@ V-B The effectiveness of InnerGCN
In this subsection, several comparison experiments are conducted to evaluate the effectiveness of InnerGCN. The hyperparameters of the model used in this section are defined in Tab.II.
@@ V-C The effectiveness of the Balancer
In fact, generative adversarial networks are subject to significant instabilities. In the models like ours, discriminators tend to reach the optimal case more quickly than generators, which leaves a great performance divide between discriminators and generators in same model.
To tackle this problem, the training speed of the discriminator need to slow down and the gap between source domain and target domain should be reduced. In other words, retarding the optimization process of the discriminator makes the generator’s training more stable.
@@ V-F Differences of brain adjacency matrices in different categories compared with NC
Based on above four comparison results, we further counted the first 30 abnormal brain connections lower than 0.05 in these experiments compared to normal subjects, and the results are shown in the Tab.III. From the table, a clear phenomenon can be found: there were multiple abnormal brain connections in the structural experiments, and these brain connections appeared many times in these four groups of experiments, while there were no such abnormal brain connections in the functional experiments, or at most in two groups of experiments. This proves that the deterioration of brain structure is gradual, but through its coordination mechanism, the brain can skip the original abnormal brain connections, use other connections to achieve the same task, which are different with the brain connections of NC.
@@
By comparing these first 30 brain connections, just shown in Tab.III, it was found that the structural data is relatively stable, such as the brain connections between brain region Occipital_Mid_R and Occipital_Inf_R or brain region Temporal_Mid_L and Temporal_Pole_Mid_L, appeared many times in these four groups of experiments. For this reason, we utilized 60 subjects and counted the connection strength of their brain connections between brain region Temporal_Mid_L and Temporal_Pole_Mid_L, and between brain region Occipital_Mid_R and Occipital_Inf_R. The statistical result is shown in Fig.7. Fig.7(a) and Fig.7(b) are the abnormal brain connections’ results, while Fig.7(c) and Fig.7(d) are the normal brain connections’ results. Abnormal brain connections exhibit more instability than normal brain connections in structure domain. Different with the brain normal connections, the curves of the abnormal connection between brain region Occipital_Mid_R and Occipital_Inf_R, or brain region Temporal_Mid_L and Temporal_Pole_Mid_L show a large degree of amplitude change. And it demonstrates that the abnormal situation got from the experiments are much different with the normal connection. The abnormal brain connections in {40-56, 85-87, 52-54, 30-72, 41-87} have the same change trend.
@@ VI Discussion
There are still some problems that are not solved in this paper. Although the correspondence between structure and function can be imitated by generators, the specific correspondence between the structure and function, i.e., the coordination mechanism of the brain, remains ungraspable. Some subjects have large differences compared to the same category of subjects, and how they should go about the diagnosis.
@@ VII Conclusion
In this paper, we proposed a new graph convolution method, called InnerGCN, to map the structural and functional brain networks of human brain to each other. And two generators and discriminators learn the structure-to-function and function-to-structure mapping function, respectively. However, direct comparison of structure and function doesn’t not reveal similarities between structure and function, suggesting that there is not a one-to-one correspondence between structure and function, but rather a complex coordination mechanism. In addition, we compared the structural and functional connectivity of the patients during the deterioration of the disease AD. The structural data yielded more stable abnormal brain connections, while the functional data yielded abnormal brain connections that were difficult to detect as such, but many of these connections were connected to some specified brain regions.

### v4

@@ Abstract
The relationship between brain structure and function is critical for revealing the pathogenesis of brain disorders, including Alzheimer’s disease (AD). However, mapping brain structure to function connections is a very challenging task. In this work, a bidirectional graph generative adversarial network (BG-GAN) is proposed to represent brain structure-function connections. Specifically, by designing a module incorporating inner graph convolution network (InnerGCN), the generators of BG-GAN can employ features of direct and indirect brain regions to learn the mapping function between the structural domain and the functional domain. Besides, a new module named Balancer is designed to counterpoise the optimization between generators and discriminators. By introducing the Balancer into BG-GAN, both the structural generator and functional generator can not only alleviate the issue of mode collapse but also learn complementarity of structural and functional features. Experimental results using Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset show that both generated structure and function connections can improve the identification accuracy of AD. The experimental findings suggest that the relationship between brain structure and function is not a complete one-to-one correspondence. They also suggest that brain structure is the basis of brain function, and the strong structural connections are majorly accompanied by strong functional connections.
@@ I Introduction
Alzheimer’s Disease (AD) has placed a huge burden on patients’ families and society as a whole. Symptoms of AD mainly include memory impairment, aphasia, apraxia, cognitive impairment, impairment of visuospatial skills, executive dysfunction, and personality and behavior changes. Although most people experience some degree of memory loss in old age, memory loss is much more severe in people with AD than in normal individuals. The term Mild Cognitive Impairment (MCI) was created to differentiate them. Some MCI patients can remain stable for years, while some subjects with AD-like brains change inevitably progress to AD. The prevalence of AD is increasing rapidly due to the fact that the proportion of the population aged 65 and over is growing faster than any other age group worldwide in recent years. Current drugs can relieve symptoms but do not prevent disease progression. Learning the relationship between brain structure and function can help doctors differentiate between AD patients and normal people (NC). Moreover, it can also help researchers to figure out the cause of the worsening condition[1, 2, 3, 4].
Recent advancements in artificial intelligence (AI)[5, 6, 7, 8] have significantly enhanced the capabilities of neuroimaging analysis, making it more reliable and sensitive than traditional cognitive assessments in detecting AD. Images of different modalities, such as structural Magnetic Resonance Imaging (sMRI), functional Magnetic Resonance Imaging (fMRI), diffusion tensor imaging (DTI) and so on, are well suited to assist physicians in diagnosing patients with the disease [9, 10, 11, 12]. Moreover, with the advantage of high temporal and spatial resolution, each of these tools can identify differences between AD patients and NC well. In the realm of AI-driven diagnostic approaches[13, 14, 15, 16], individual modality is used to study diagnosis between AD and MCI. Some used sMRI brain neuroimages to classify AD and NC [17, 18, 19, 20, 21, 22]. Some used fMRI data [23] and some used DTI data [24]. Due to the limitations of computing power and algorithm at that time, they couldn’t utilize the complementary information between multiple modalities’ data[25]. The complex relationships between brain structure and function could not be discovered clearly. The results of diagnosis are not so good as the current methods using multiple modalities images. Now, with the proliferation of artificial intelligence technologies, it is possible to integrate data from multiple imaging modalities.
@@ II Related Works
Zhang et al. [28] recovered the structural connectivity via Multi-GCN from functional connectivity. The recovered result that functional information can predict structural information, revealed that functional information contains all the structural information. Wang et al. [29] proposed an algorithm called CDA and the empirical structural connections (SC) can be reliably predicted based on the direct anatomical relationships, indirect pathways, and module topology interacting with one another forming temporally dependent functional connections (FC). Based on the fact that related brain regions with strong functional connectivity can exist without structural connectivity, but the strong functional connectivity can be observed in the brain regions with the strong structural connectivity, there is a many-to-one functional-structural mode.
@@
However, not all researchers agree with the above views. Abdelnour et al.[30] tried to use mathematical methods to derive a relationship between SC from DTI and FC from fMRI and found that using the approach via (structural) Laplacian spectral, FC and SC shared eigenvectors and their eigenvalues are exponentially related. Honey et al.[31] believed that although resting state functional connectivity is variable, its strength, persistence and spatial statistics are constrained by the large-scale anatomical structure. Messé et al.[32] explored the spatial consistency of the relationship between the structure and function of the human brain. The result showed that brain regions can affect the relationship between structural and functional connections. Different division methods will affect the relationship between structure and function. Simon et al.[33] discussed the influence of different brain regions on structure and function and proved its effectiveness. Koch et al. [34] examined the correlation of spontaneous fluctuations of brain voxel BOLD signals in the cortical gyrus of the same hemisphere and found that there is a positive correlation between function and structure. Annen et al. [35] explored the relationship of function fluorodeoxyglucose FDG-PET metabolism and structure MRI-diffusion-weighted images (DWI). The fact that the patients’ local metabolism and white matter integrity were significantly reduced, showed that a stronger relationship of function-structure exists in most areas of Default mode networ. Straathof et al. [36] combined multiple articles and draw a conclusion: the structural and functional connection strength of the mammalian brain is positively correlated with the connection strength of the diffuse structure on the macroscopic scale, and positively correlated with the connection strength of the neuron trace structure on the mesoscale.
@@
Based on above analysis, the following simple assumption can be summarized: there is indeed a relationship between structure and function. However, very few studies focus on the relationship between structure and function by employing complementary features between brain structure and function.
@@ IV-B BGGAN for mapping domain
Generative adversarial networks [37, 38, 39, 40, 41] commonly consist of two modules: the module generator and the module discriminator. The generator tries to map latent features $z\sim p(z)$ from the noise or from the prior distribution to the distribution of the real domain. The discriminator aims to distinguish between the real data and the fake data. They are trained by the following min-max optimization Eq. 1:
@@
In addition, because of the instability of the traditional GAN framework and the unsmoothness of the training process, the results of generators are very poor. The generated data are quite different with the source data. In view of above problems, the Balancer modules are add between the source data and the real input of discriminators, which can be seen in section IV-D in details. The loss function of the model can not only calculate the differences in brain region level, but also take into account the overall differences. The details of Balancer modules can be seen in section IV-E.
@@ IV-C InnerGCN for extracting graphic features
Next, we will introduce how to get $U$ in next steps: 1) get the Laplace tensor $L$, where $L(i)=I-D(i)^{-1/2}A(i)D(i)^{-1/2}$; 2) get the $L_{dft}$ from the $L$ in the Fourier form; 3) get the singular values matrix of each dimension of the tensor $[u(i),s(i)]=EVD(L_{dft}(i))$; 4) map the data in the Fourier data back to the original space $U=IDFT(u),\Lambda=IDFT(s)$. $U^{H}$ is the Hermitian matrix of $U$. $U\diamond x$ is the $U$ of Fourier form, and same as convolution filter $g$. The third-order tensor can be considered as matrix with each element to be a vector. Same with matrix multiplication based on vector multiplication, matrix-matrix interaction is realized through circular convolution based on matrix multiplication. We defined $A,B$ as $A\in\mathbb{R}^{n_{1}\times n_{2}\times 3}$ and $B\in\mathbb{R}^{n_{2}\times n_{3}\times 3}$ and fold1(…), fold2(…), unfold(…) in Eq (6-9):
@@ IV-D Balancer for smoothing training process of BGGAN
A neural network is utilized as Balancer, which is comprised of several convolution networks, a step connection combining the high order information and the low order information from the Balancer. Its loss function contains two parts: an adversarial loss $\mathcal{L}_{B}^{adv}$ and a reconstruction loss $\mathcal{L}_{B}^{recon}$. The function of the adversarial loss is to try fool the discriminator and slow down the discriminators’ learning process. The reconstruction loss is to limit the output similar to real sample and avoid the discriminator learn incorrect distribution too much. The Balancer is trained as following loss:
@@ V Experiments
V-A Datasets and Experimental Settings
In this subsection, a series of experiments are conducted based on the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [48]. In our experiments, we used 80% of the subjects as the training dataset and 20% of the subjects as the testing dataset. All experimental results are based on the testing dataset. The details of the datasets used in this paper are presented in Tab.I, which provides the number of subjects, gender distribution, mean age, and mean weight (with standard deviations) across five categories: Normal Control (NC), Significant Memory Concern (SMC), Early Mild Cognitive Impairment (EMCI), Late Mild Cognitive Impairment (LMCI), and Alzheimer’s Disease (AD). This detailed summary allows for a clear understanding of the dataset composition and demographic distribution used in the experiments.
@@ V-B The effectiveness of InnerGCN
In this subsection, several comparison experiments are conducted to evaluate the effectiveness of InnerGCN. The hyperparameters of the model used in this section are defined in Tab.II, which outlines the optimal parameter settings, including the optimizer type, learning rates for generators and discriminators, batch size, stop strategy, number of regions of interest (ROIs), and the modalities used (sMRI, DTI, and fMRI). These parameters were selected based on preliminary experiments and aim to ensure the most robust and accurate model performance in the conducted experiments.
@@ V-C The effectiveness of the Balancer
In fact, generative adversarial networks are subject to significant instabilities. In the models like ours, discriminators tend to reach the optimal case more quickly than generators, which leaves a great performance divide between discriminators and generators in the same model.
To tackle this problem, the training speed of the discriminator needs to slow down and the gap between source domain and target domain should be reduced. In other words, retarding the optimization process of the discriminator makes the generator’s training more stable.
@@ V-F Differences of brain adjacency matrices in different categories compared with NC
Based on above four comparison results, we further counted the first 30 abnormal brain connections lower than 0.05 in these experiments compared to normal subjects, and the results are shown in the Tab.III. From the table, a clear phenomenon can be found: there were multiple abnormal brain connections in the structural experiments, and these brain connections appeared many times in these four groups of experiments, while there were no such abnormal brain connections in the functional experiments, or at most in two groups of experiments. This proves that the deterioration of brain structure is gradual, but through its coordination mechanism, the brain can skip the original abnormal brain connections, use other connections to achieve the same task, which are different from the brain connections of NC.
@@
By comparing these first 30 brain connections, just shown in Tab.III, it was found that the structural data is relatively stable, such as the brain connections between brain region Occipital_Mid_R and Occipital_Inf_R or brain region Temporal_Mid_L and Temporal_Pole_Mid_L, appeared many times in these four groups of experiments. For this reason, we utilized 60 subjects and counted the connection strength of their brain connections between brain region Temporal_Mid_L and Temporal_Pole_Mid_L, and between brain region Occipital_Mid_R and Occipital_Inf_R. The statistical result is shown in Fig.7. Fig.7(a) and Fig.7(b) are the abnormal brain connections’ results, while Fig.7(c) and Fig.7(d) are the normal brain connections’ results. Abnormal brain connections exhibit more instability than normal brain connections in the structural domain. Different with the brain normal connections, the curves of the abnormal connection between brain region Occipital_Mid_R and Occipital_Inf_R, or brain region Temporal_Mid_L and Temporal_Pole_Mid_L show a large degree of amplitude change. And it demonstrates that the abnormal situation got from the experiments are much different with the normal connection. The abnormal brain connections in {40-56, 85-87, 52-54, 30-72, 41-87} have the same change trend.
@@ VI Discussion
This work hasn’t been able to solve all the prevailing issues related to brain structure and connection mapping. Although the correspondence between structure and function can be imitated by generators, the specific correspondence between the structure and function, i.e., the coordination mechanism of the brain, remains ungraspable. Some subjects have large differences compared to the same category of subjects, and how they should go about the diagnosis.
@@ VII Conclusion
In this paper, we proposed a novel graph convolution method, called InnerGCN, designed to map the structural and functional brain networks of the human brain to one another. The method utilizes two distinct generators and discriminators to learn the structure-to-function and function-to-structure mapping functions, respectively. Our findings indicate that a direct comparison between structural and functional networks does not reveal straightforward similarities, suggesting the absence of a one-to-one correspondence between structure and function. Instead, this highlights the presence of a complex coordination mechanism underlying their interactions. Furthermore, we analyzed the structural and functional connectivity of patients experiencing Alzheimer’s Disease (AD) progression. The results show that structural data revealed more stable abnormal brain connections, whereas functional data identified abnormal connections that were less apparent but often associated with specific brain regions. These findings suggest that structural and functional networks provide complementary insights into brain connectivity during disease progression. Future work will explore more advanced coordination models to better capture the complex interactions between structural and functional networks, as well as investigate the generalizability of InnerGCN across other neurodegenerative diseases.
Acknowledgment
This work was supported by the National Natural Science Foundations of China under Grant 12326614 and 62172403, the Distinguished Young Scholars Fund of Guangdong under Grant 2021B1515020019.